How many liters of a 30% alcohol solution must be mixed with 80 liters of a 90% solution to get an 80% solution?
Answer by checkley77(12844)

.90*80+.30x=.80(80+x)
72+.30x=64+.80x
.30x-.80x=64-72
-.50x=-8
x=-8/-.50
x=16 liters of 30% alcohol is used.
Proof:
.90*80+.30*16=.80(80+16)
72+4.8=.8*96
76.8=76.8